find the mass and center of mass of the triangular lamina with vertices (0,0), (0,1) and (1,0) and density function  𝛿(x,y)=xy


1
Expert's answer
2020-07-14T18:23:23-0400

Let "A=(0,1), B=(1, 0), O=(0,0)."

Line OA: "x=0, 0\leq y\leq1"

LineAB: "y=-x+1, 0\leq x\leq1"

Line OB: "y=0, 0\leq x\leq 1"



Given "\delta(x,y)=xy"

Find the mass of the lamina


"m=\iint_D\delta(x,y)dA=\displaystyle\int_{0}^1\displaystyle\int_{0}^{1-x}xydydx="

"=\displaystyle\int_{0}^1x\big[{y^2\over 2}\big]\begin{matrix}
 1-x \\
 0 
\end{matrix}dx={1\over 2}\displaystyle\int_{0}^1(x-2x^2+x^3 )dx="

"=\big[{x^2 \over 4}-{x^3 \over 3}+{x^4 \over 8}\big]\begin{matrix}
 1 \\
 0
\end{matrix}={1 \over 24} (units\ of\ mass)"

Mass of the lamina is "\dfrac{1}{24}" units of mas.


Find the coordinates of the center of mass


"\bar{x}={1\over m}\iint_Dx\delta(x,y)dA=24\displaystyle\int_{0}^1\displaystyle\int_{0}^{1-x}x^2ydydx="

"=24\displaystyle\int_{0}^1x^2\big[{y^2\over 2}\big]\begin{matrix}
 1-x \\
 0 
\end{matrix}dx=12\displaystyle\int_{0}^1(x^2-2x^3+x^4 )dx="

"=12\big[{x^3 \over 3}-{2x^4 \over 4}+{x^5 \over 5}\big]\begin{matrix}
 1 \\
 0
\end{matrix}=4-6+{12 \over5}={2 \over5}"

"\bar{y}={1\over m}\iint_Dy\delta(x,y)dA=24\displaystyle\int_{0}^1\displaystyle\int_{0}^{1-x}xy^2dydx="

"=24\displaystyle\int_{0}^1x\big[{y^3\over 3}\big]\begin{matrix}
 1-x \\
 0 
\end{matrix}dx=8\displaystyle\int_{0}^1(x-3x^2+3x^3-x^4 )dx="

"=8\big[{x^2 \over 2}-{3x^3 \over 3}+{3x^4 \over 4}-{x^5 \over 5}\big]\begin{matrix}
 1 \\
 0
\end{matrix}=4-8+6-{8 \over5}={2 \over5}"

Center of mass "\big(\dfrac{2}{5},\dfrac{2}{5}\big)"